OpenStudy (rizags):

Intern E.T. phoned home for some monetary help from his parents. He asked them for 23144 Asogi dollars to pay his rent. Asogi numbers are in base 6 (E.T. had only 3 fingers on each hand), and 7 Asogi dollars are equivalent to 1 American dollar. How many American dollars did E.T. pay for rent?

OpenStudy (aum):

Convert 23144 in base 6 to decimals. Divide it by 7 to convert it to dollars.

OpenStudy (rizags):

how do i convert between bases? is it like modular arithmetic where a normal 10 would turn to a 3?

OpenStudy (aum):

Look at how the decimal number system works: 362 in decimal = 2 + 6*10 + 3*10^2 = 2 + 60 + 300 = 362 23144 in base 6 = 4 + 4*6 + 1*6^2 + 3*6^3 + 2*6^4 = ?

OpenStudy (rizags):

oh i get it. its 3304 in base 6

OpenStudy (rizags):

but wait isnt it already in base 6?

OpenStudy (aum):

Yes. 23144 in base 6 = 3304 in base 10 (or in decimals).

OpenStudy (rizags):

how would i convert base 10 into base 6?

OpenStudy (aum):

Now we have a decimal number. Divide it by 7 to convert to dollars and express the result in decimals so normal humans can understand.

OpenStudy (rizags):

well it is 472 dollars and thank you very much but how would base 10 be changed to base 6?

OpenStudy (aum):

To convert decimal number to base 6 you have to keep dividing by 6. Convert 3304 in decimal to a base 6 number: 6 | 3304 --------- 6 | 550 R:4 |------- 6 | 91 R: 4 -------- 6 | 15 R: 1 -------- 2 R: 3 23144

OpenStudy (rizags):

the remainders make up the number because they are what is left when 3304 is divided by all the powers of 6 included in the original 3304 correct?

OpenStudy (rizags):

i mean the original 23144

OpenStudy (aum):

correct. in the original 3304.

OpenStudy (rizags):

ok thank you so much i appreciate it

OpenStudy (aum):

you are welcome.